Soft drinks are commonly sold in aluminum containers. How many such containers are thrown away or recycled each year by consumers ? (Assume an average of 3 cans per person each week and a population of 320 million Enter your answer using one of the following formats: 1.20e- 3 for 0.0012 and 1.20e+2 for 120. How many tons of aluminum does this represent?(assume an average weight of 0.50 oz of aluminum per can) 


1
Expert's answer
2021-02-04T02:32:38-0500

There are approximately 51 weeks in year. Thus, each person throw away "51\cdot 3 = 153" cans per year. The population of 320 million uses



"153\cdot 320\times 10^6 \approx 4.9\times10^{10}"

cans per year.

Since the weight of each can is "0.5\space oz \approx 0.014\space kg", the total weight is:


"0.014\cdot 4.9\times 10^{10} = 6.86\times 10^8 kg = 6.86\times 10^5tons"

Answer. 4.9e+10 cans, 6.86e+5 tons.